Crew 335 HSO Beginning Of Mission Checklist 19Apr2026

[title HSO Beginning Of Mission Checklist- April 19th]

Report title: HSO Beginning Of Mission Checklist

Crew #: 335

Position: HSO

Report prepared by: MATTEO CALORE

Date: 19-04-2026

Part 1

Locate and confirm the emergency escape routes in the Hab are functional and clear:

Stairs (between lower end upper deck) – stairs between L and U decks
Emergency window (upper deck, east side) – emergency window on U deck
Commander’s window (located in the commander’s crew quarter)

All the emergency routes are functional and clear
Part 2

Inventory First Aid kit

BIG ONE WITH MIRROR:
EMERGENCY BLANKET X5
SPLINT RHINO RESCUE X1
TRIANGULAR BENDAGE X3
ELASTIC BENDAGE X2
DRAMAMINE TABLETS X1 (expired on the 09/2025)
ADESIVE TAPE X3 ROLLS
SELF-ADHERING SPORTS WRAP X1
ALCOHOL PREP PADS
COTTON SWABS X400
ERLOOP FACE MASK X19
IBUPROFEN – almost finished ONLY X2 200MG
MELANGE FOR NASAL CONGESTION X1
ADHESIVE BENDAGE X19
ELETRONIC BLOOD PRESSION MONITOR 1X WITH BATTERIES
FINGERTIOPS PULSE OXIMETER X1
INFRARED THERMOMETER X1
ISOPROPYL ALCOHOL X1 BOTTLE
HYDROGEN PEROXIDE TOPIC X1 BOTTLE

SMALL ONE AMERICAN RED CROSS:

EYEWASH X1 // EXPIRED 2024
ANTIBIOTIC NEOMYCIN almost finished ONLY ONE SMALL PACKET // EXPIRED 2025
BURN DRESSING X1 PACKET
GAUZE X 4
SCISSOR X 1
TWEEZERS X1

Needs to be refilled:

Inventory First Aid kit and note what needs to be refilled:
DRAMAMINE TABLETS
IBUPROFEN
ANTIBIOTIC
FOILLE FIRST AID OINTMENT

Part 3

Note any safety issues: none

Note any health/environmental issues: none

Note any missing or recommended health and safety supplies: a defibrillator

Part 4. Using the attached Safety Equipment Inventory, locate, test and confirm operation of all safety equipment. List any equipment not found and/or missing. See notes on the next page.

Safety Equipment Inventory

HAB Upper deck HAB Lower deck RAM Green Hab Science Dome
Escape ladder X N/A N/A N/A N/A
Eyewash N/A N/A N/A N/A X
Fire blanket X N/A X N/A X
Fire extinguisher X X X X X
First Aid N/A X N/A N/A X
Intercom X N/A X X X
Radios (Channels 10 and 22) X X X X X
Nightlight X X X N/A X
Carbon Monoxide alarm X X X X X
Smoke alarm X X X X X
Propane alarm X X N/A N/A N/A
EVA Safety Kit N/A X N/A N/A N/A

Journalist Report – April 25th

Crew #: 335

Position: Crew Journalist

Report prepared by: MatteoCalore

Date: 25APR2026

Sol: 6

Report Title: Returning to Form

This morning, a few of us woke to a rather unusual kind of alarm. A strange little creature slipped into our sleeping quarters, calling us back to the land of the awake. A surreal and slightly disorienting way to begin the day.

Later, three of us set out on a long morning EVA to Somerville Overlook. As always, the landscape was breathtaking. Dave focused on capturing a panoramic sequence, while the filmmakers worked on documenting the terrain, training their gaze and cameras to hold the vastness of these endless red valleys. It became another productive and rewarding day of filming.

We spent the rest of the day in the science dome, working to activate the microwave kiln for Charlotte’s clay project. It required patience, along with several trials and adjustments, but eventually everything came together. We finally witnessed the first clay experiment take form—our “sixth crewmate,” the small desert mouse we encountered yesterday. After about twelve minutes of firing in the microwave and nearly an hour of cooling, the mouse was preserved in the earth collected from the canyons. A small but meaningful transformation—matter shifting state, holding memory.

Our crew engineer also had a very full day. She dedicated significant time to processing Messier 16 astronomy data and continued her rover research.

We ended the day sharing a warm meal and looking at photographs of Fremont culture clay artifacts taken by our commander. A day marked by cycles of observation, transformation, and quiet reconnection.

Supplemental Operations Report – April 25th

Date: 4/25/2026
Name of person filing report: Ben Stanley
Reason for Report: Routine
Non-Nominal Systems: Hab outer shell. Director’s trailer roof.

Power system:
1) Power system: Nominal.
2) Current generator hours: 1449.0

Propane Readings:
Station Tank: 46%
Director Tank: 56%
Intern Tank: 70%
Generator Tank: 80%

Water:
Hab Static Tank – 161 gallons above safety margin
GreenHab – 91 gallons above safety margin
Outpost tank – 460 gallons
Septic Tanks emptied: No

Rovers:
Sojourner rover used: No
Hours: 226.9
Beginning Charge: 100%
Ending Charge: 100%
Currently Charging: Yes

Cars:
Hab Car used and why, where: Multiple trips to Hanksville for supplies
Crew Car used and why, where: Replacement vehicle is still being sourced
General notes and comments: The station will be operating with the hab car only until crew car function is replaced

Summary of Internet: Nominal.

EVA suits and radios:
Suits: Nominal.
Comms: Crew reports some radios are malfunctioning.
T-Echo EVA-link: Nominal

Campus wide inspection, if action taken, what and why: Hab outer shell shows ongoing wear.
Summary of Hab Operations: Nominal.
Summary of GreenHab Operations: Nominal.
Summary of SciDome Operations: Nominal.
Summary of Observatories Operations: Nominal.
Summary of RAM Operations: Nominal.
Summary of Outpost Operations: Roof leak remains unaddressed, but have not had rain. Will execute repair this week.
Summary of Health and Safety Issues: Some crew members have been reporting mild to moderate cold- or flu-like symptoms.

Journalist Report – April 30th

Journalist Report

Crew #: 335

Position: Matteo Calore (Crew Journalist)

Date: 30APR2026

Sol: 11

Today is our last day on Mars. We woke up around 7:00 a.m. and started the day with pancakes. We were all a bit tired, but excited for our final EVA. We were finally heading to the Sea of Shells, and we were all curious to see this new territory. We have to say that reality exceeded our expectations.

When we arrived at the Mars Overlook, we were all amazed by the beauty and diversity of the landscape. The red gradients of Mars slowly faded from a light beige-brown into a dark grey and black, transporting us into a lunar landscape. We were no longer on Mars, but on the Moon, or perhaps on one of Mars’ moons.

We paused to take in this extraordinary view before slowly descending toward the Sea of Shells. There, we found ourselves walking on a carpet of shells of indeterminate age, mysterious fossils from a time unknown to us. We wondered what kind of shells they were and how they could have ended up there in such vast quantities. These questions remained unanswered. In the distance, we could see a black obelisk, almost like a spacecraft ready to launch into space.

We then made our way back slowly, taking in the Martian landscape one last time through our suits. As our Commander says, it is difficult to stop being amazed by the beauty, diversity, and sense of wonder that these places evoke.

Back at the Hub, we shared our final meal as astronauts. This afternoon, we ended the simulation and returned to Earth. It is a strange feeling to shift from one planet to another. To know that now you are surrounded by a civilization, by shops, by restaurants, and that you can simply get into a car and go wherever you want.

EVA Report – April 30th

EVA Report

Crew #: 335

Position: Elisa Strinna (Crew Artist)

EVA #: 14

Date: 30APR2026

Sol: 11

Purpose of EVA: TTo visit the Sea of Shells for film shooting. Crew members Elisa Strinna and Matteo Calore will focus on filming activities. Charlotte Brugge will be filmed while exploring the site in search of traces of past life. Commander David Laude will conduct a panoramic photographic survey at one or two locations.

Weather Conditions: Overall good conditions, sunny with temperatures between 53°-64° F

Start Time: 10:15

End Time: 13:05

Narrative: Around 10:15, we headed toward the Sea of Shells. The ride was very pleasant, as the landscape shifted through different gradients and textures. We stopped at the Grey Overlook to admire the panorama for about half an hour, and then continued down to the Sea of Shells.

We spent another half hour observing the fossil formations before beginning our return to the Hab. On the way back, we made a final short stop to take in the Mars-like landscape one last time. The film crew documented the entire EVA, while the Commander focused on his project of capturing panoramic views of the different locations.

Destination (UTM WGS84): Sea of Shells (4255500N, 516500E)

Road(s) and Routes (per MDRS Map): Cow Dung Road 0110 → Brahe Hwy 1572/1575 → Copernicus Hwy 1575 → final approach on foot to Sea of Shells.

Mode of Travel: Rover + Walking

EVA Participants: David Laude (Commander) Elisa Strinna (Crew Artist), Matteo Calore (Crew HSO / Journalist) Charlotte Brugge (GreenHub Officer)

Operations Report – April 30th

Report title: Operations Report

Crew #: 335

Position: Commander

Report prepared by: Dave Laude

Date: 30Apr2026

Sol: 11

Non-nominal systems: None

Notes on non-nominal systems: None

ROVERS

Spirit rover used:

Hours: 311

Beginning charge: 100

Ending charge: 40

Currently charging: Yes

Opportunity rover:

Hours: 274.4

Beginning charge: 100

Ending charge: 34

Currently charging: Yes

Curiosity rover used:

Hours: 0

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Perseverance rover used:

Hours: 0

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

General notes on rovers: Some key switches may need lubricant, especially Opportunity.

Summary of Hab operations:

Water Use (see notes): 30 gal/d

Main tank level (remaining gallons, see notes): 136 gal

Main water tank pipe heater (ON or OFF): OFF

Main water tank heater (ON or OFF) OFF (was removed)

Toilet tank emptied (NO or YES): NO

Summary of internet: Connection was used for research, accessing, and working on homework, sending reports, video editing and listening to music.

Summary of suits and radios: EVA suit #7 non-functional as described yesterday.

Summary of GreenHab operations: (use information from the GreenHab report)

Water use: 1 gallon. Currently being dismantled.

Heater (ON or OFF): Auto

Supplemental light (hours of operation): Auto

Harvest (name, weight in grams): 0g

Summary of ScienceDome operations: Used to film creative projects, drying clay and manipulating clay

Dual split (Heat or AC, On or Off or Automatic): Automatic

Summary of RAM operations: Not used.

Summary of any observatory issues: Not used.

Summary of health and safety issues: None

Questions, concerns and requests to Mission Support: None

Sol Summary – April 30th

Report Title: Sol Summary Report

Crew #: 335
Position: Crew Artist
Report Prepared by: Elisa Strinna
Date: 30APR2026
Sol: 11

Summary Title: Moon Overlook

Mission Status: We are at the end of our simulation, the mission was successful.

This morning we woke up at 7. Some crew members made pancakes, which were quite delicious. After breakfast, we began preparing for the EVA, getting dressed and organizing the equipment. Around 10:15, we headed toward the Sea of Shells. The ride was very pleasant, as the landscape shifted through different gradients and textures. We stopped at the Moon Overlook to admire the panorama for about half an hour, and then continued down to the Sea of Shells.We spent another half hour observing the fossil formations before beginning our return to the Hab. On the way back, we made a final short stop to take in the Mars-like landscape one last time. The film crew documented the entire EVA, while the Commander focused on his project of capturing panoramic views of the different locations.

Once back at the Hab, we had a satisfying lunch of bean balls and mashed potatoes, along with cabbage fresh green leaves from the greenhouse. The Commander surprised us with cherry-chocolate bonbons for dessert.

After lunch, we dismantled the greenhouse, and around 16:30 we ended the simulation, finally returning to Earth.

Anomalies in Work: We decided to end SIM earlier because the film crew needs to film some locations around MDRS out of SIM.

Weather: Conditions were ideal with temperatures of 68F and clear sky.

Crew Physical Status: Most crew members are generally in good condition. A little bit of tiredness.

Sol Summary – April 29th

Report Title: Sol Summary Report
Crew #: 335
Position: HSO
Report Prepared by: Matteo Calore
Date: 29 Apr 2026
Sol: 10

Summary Title: Walking on Mars

Mission Status: Mostly in good shape with continuing daily progress.

Sol Activity Summary: Today we woke up as usual between 7 and 8. The day began in a dynamic way, with many discussions about the development of our projects. Afterwards, the film crew continued conducting interviews with the other crew members. The interviews went very well, offering many interesting insights into personal lives and the motivations that brought us here, as well as what draws people toward space science and the mystery that surrounds the universe.

In the afternoon, we continued working on our projects. As we are approaching the end of our simulation, we are mostly focusing on wrapping things up, trying to understand what might still be missing and what can still be done. Toward evening we took a walk on the highest hills behind the MDRS. The light was particularly beautiful, as was the view of the desert from our habitat. The gradual transitions of color across the clay hills—from red to white, to violet, to beige—and the scattered rocks that populate them make this experience truly unique.

For both lunch and dinner, the Italians prepared two different types of pasta: the first with a tuna ragù, and the second a kind of baked macaroni and cheese.

Anomalies in Work: EVA was canceled due to unforeseen circumstances.

Weather: Conditions were ideal with high of 68F, light winds and slight cloud cover.

Crew Physical Status: Most crew members are generally in good condition. A couple are experiencing mild symptoms such as sore throats and nasal congestion.

Operations Report – April 29th

Report title: Operations Report

Crew #: 335

Position: Commander

Report prepared by: Dave Laude

Date: 29Apr2026

Sol: 10

Non-nominal systems: None

Notes on non-nominal systems: None

ROVERS

Spirit rover used:

Hours: 0

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Opportunity rover:

Hours: 0

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Curiosity rover used:

Hours: 0

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Perseverance rover used:

Hours: 0

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

General notes on rovers: Some key switches may need lubricant, especially Opportunity.

Summary of Hab operations:

Water Use (see notes): 32.94 gal/d

Main tank level (remaining gallons, see notes): 156 gal

Main water tank pipe heater (ON or OFF): OFF

Main water tank heater (ON or OFF) OFF (was removed)

Toilet tank emptied (NO or YES): NO

Summary of internet: Connection was used for research, accessing, and working on homework, sending reports, video editing and listening to music.

Summary of suits and radios: EVA suit #7, the single piece suit, had a ventilation failure near the end of sim yesterday. The fan was found non-functional. Dave opened the back and immediately saw a thick red wire dangling near the battery. This was the + battery connection. There was no reason for this found and remains a mystery. The connection was remade and found to be sound. It appears that any force on the wire due to its position, weight and bending would apply force downward thus reinforcing the connection. Once reattached, there was no fan function so the lower access panel was opened to troubleshoot more and one switch terminal to the fan was loose. Not just the wire held by screw was loose, but also the switch lug was loose in its molded plastic. Wiggling that connection occasionally powered up the fan. A replacement switch was not found with the replacement parts. Please have Scott Davis advise. This suit is not needed now.

Summary of GreenHab operations: (use information from the GreenHab report)

Water use: 6 gallons

Heater (ON or OFF): Auto

Supplemental light (hours of operation): Auto

Harvest (name, weight in grams): 0g

Summary of ScienceDome operations: Used to film creative projects, drying clay and manipulating clay

Dual split (Heat or AC, On or Off or Automatic): Automatic

Summary of RAM operations: DVM and 2 screw drivers removed to troubleshoot EVA suit #7. Will replace tools soon.

Summary of any observatory issues: Not used.

Summary of health and safety issues: redacted.

Questions, concerns and requests to Mission Support: None